Solve P = 1/6 (r + 5t) for r. You must show all work.
Found 2 solutions by Fombitz, edjones:
Answer by Fombitz(32388)
Answer by edjones(8007)

P = 1/6 (r + 5t)
6p=r+5t multiply 6 times each side.
6p-5t=r subtract 5t from each side
.
Ed